A Bibliography of Publications of Luc Steels

Luc Steels

Tel: ?n/a? FAX: ?n/a?

E-mail: elisahus@vub.ac.be (Internet)

13 October 2017 Version 1.13

Abstract

This bibliography records publications of Luc Steels.

Title word cross-reference

1986 [dBHS87]. 1999 [PVC^{+}99].

4 [Ste76a].

5th [FMN99].

6th [SDR^{+}00].

8th [SSv94b, SSV94a].

’94 [SSv94b, SSV94a]. 97 [vW97]. ’99 [PVC^{+}99]. 9th [vW97].

Acquisition [BSV98, SSV94a, SBV98, SSv94b, Ste94f, Ste97d, Ste02f]. Action [Ste00e]. adaptatifs [SV97a]. Adaptive [C^{+}94, MMM^{+}96, MW90, MBF^{+}00, PBMW98, Ste96c, Ste98d, Ste99c, Ste00d, SV97b]. Advances [FMN99, dBHS87]. age [Ste97c]. Agent [Ste98b]. Agents [Ste94d, SB95, Ste95b, Ste95k, Ste95a, Ste97g, Ste00a, Ste90b, Ste90c, Ste94c, Ste94h, Ste95j, SSV96, SV97a, SV97b, Ste98c]. AI [Ste84c, Ste86a, Ste89a]. AIBO [SK01]. AISB91 [SS91]. algorithm [dJS03]. Algorithmic [Ste85a]. Amorcage [SK99a]. Analogical [Ste90e, Ste90d]. Analysis [Ste94e, Ste94g, PVC^{+}99]. Ancrage [SV97a]. ancrés [SK99a]. animals [C^{+}94, MMM^{+}96, MW90, MBF^{+}00, PBMW98]. animats [C^{+}94, MMM^{+}96, MW90, MBF^{+}00, PBMW98]. Application [Ste90g]. applications [Ste75a]. applicative [Ste83b]. April [vW97, SS91]. Architecture [KKS95]. architectures [ADS^{+}87]. Ars [GW95]. art [NKR^{+}85]. Artefacte [Ste00c]. article [Ste01d]. Artificial [IPSK04, SC85, Ste87a, Ste88a, Ste88b, SS91, Ste91b, Ste94a, SB95, ST95, Ste95h, Ste01d, FMN99, Ste95g, dBHS87, Ste95c]. Artificial-Intelligence-Systemen [Ste95c]. Artificielle [Ste01c]. Aspects [Ste77]. automata [SV76]. autonomes [SK99a]. Autonomous [Ste93d, Ste95b, Ste95k, Ste95a, SdB95, Ste00a, Ste01b, Ste94c, Ste94h, Ste95j, SSV96, Ste97d].

Barbara [Ste01d]. based [SS83]. Basic [SdB95]. be [Ste01a]. Behavior [C^{+}94, KKS95, MMM^{+}96, MW90, MBF^{+}00, PBMW98, Ste94e, Ste94g, Ste95b, Ste94c, Ste97d]. behavior-oriented [Ste94c]. Behavioral [BKS00]. Behaviour [SS91]. Behaviours [SdB95]. Belgium [SSv94b]. better [Ste02d]. between [Ste90b, Ste90c]. Beyond [Ste94b]. Biology [Ste95k, Ste95a]. Bootstrapping [SK02]. Bootstraps [Ste99b]. Brighton [dBHS87]. Building [SB95, Ste95b, GSV93, Ste84c].

Cambridge [Ste90g]. Can [Ste01a, Ste94f]. Case [Ste76b, Ste94c, Ste02a]. Castle [IPSK04]. Categories [BSV98, SBV98, Ste98a, KS00]. challenges [Ste94h]. Change [SK98a]. channels [JS99]. choice [Ste91a]. Churchill [Ste90g]. Co [Ste98e]. Co-Evolution [Ste98e]. Cognition [Ste99b]. Cognitive [Ste98d, Ste99a]. Collectieve [Ste86b]. Collective [SK99b]. College [Ste90g]. Comment [KS00]. Commentary [SM94, Ste01d]. Commet [GSV93]. Common [Ste88d]. Communicating [Ste00a, Ste79b]. Communication [Ste98b, Ste01f, Ste97c, Ste02e, dJS03]. Community [Ste86c]. Competitors [Ste96b]. Completion [Ste75a, SV76]. Complex [Ste88a, Ste88b, Ste00d]. Componential [Ste93a]. Components [Ste90a]. comprendre [Ste02d]. computational [Ste76c]. Computer [Ste02a, van95]. Computing [ST95]. Concepts [Ste85a]. Conference [C^{+}94, FMN99, MMM^{+}96, MW90, MBF^{+}00, PBMW98, SDR^{+}00, vW97, Ste90g, SS91, dBHS87, Ste76c]. Connectionism [PSFSS89, PSFSS91]. Connectionist [Ste89a]. conscious [Ste01a]. consciousness [Ste95g]. Constraints [Ste82, Ste83a, Ste85b, SO00]. Constructing [Ste97a]. Construction [BSV98, SBV98]. construisent [KS00]. Consultants [Ste82, Ste85b]. Contact [Ste97b]. context [Ste75b]. context-free [Ste75b]. Contributions [SL92]. Conventions [Ste98b]. convergence [KS00]. Conversations [SM94]. Cooperation [Ste90b, Ste90c]. Corporate [Ste93b, Ste93c]. Coupling [Ste98d]. Creation [Ste96e]. Creatures [Ste91b]. Crucial [SKMV02]. cultural [SO00, Ste02e]. cyber [Ste95e]. cyber-sapiens [Ste95e]. cyberspace [Ste97c]. Czech [vW97].

d’agents [SK99a]. Dagstuhl [IPSK04]. dans [SV97a, SK99a]. Deepening [Ste88c, Ste89b]. delivery [Ste91a]. description [ADS^{+}87]. Descriptions [Ste83a]. Design [Ste85c, Ste94c]. development [Ste91a, dJS03]. developments [van95]. Diagnosis [Ste89c]. Discovering [Ste96b]. Discussion [Ste91a]. Distinctions [Ste97a]. Distributed [SM99, Ste90b, Ste90c, dJS03]. d’une [SK99a]. Dynamics [Ste88a, Ste88b, Ste95f, SK99b].

ECAI [dBHS87]. ECAI-86 [dBHS87]. ECAL’99 [FMN99]. ECML [vW97]. ECML-97 [vW97]. editorial [Ste95a]. Efficient [BKS00]. eighth [SS91]. EKAW [SSV94a, SSv94b]. Electronica [GW95]. Embodied [SB95, IPSK04]. Embodiment [Ste99a]. Emergence [Ste00a]. Emergent [Ste90f, Ste91b, Ste94d, Ste96c, Ste02c]. Engeln [Ste00b]. Engineering [SL92, SL93]. Enhancing [SL92]. Equilibrium [Ste94e]. erfinden [Ste00c]. erwachen [Ste95e]. ESPRIT [SL92, SL93]. EuroEcTel [PVC^{+}99]. Europal [Ste90g]. European [FMN99, PVC^{+}99, vW97, Ste90g, SSv94b, SSV94a, dBHS87]. EvoIASP [PVC^{+}99]. Evolution [Ste94d, Ste98e, SO00, Ste00f, Ste02e, SK00]. Evolutionary [PVC^{+}99, Ste02b].

Evolutionnaire [Ste01c]. examples [SS83]. Expériences [KS00, SK00]. Experiment [Ste99d, Ste01g]. Expert [Ste84a, Ste85d, Ste85e, Ste87d, Ste88c, Ste89b, SV89, SM94, Ste86c, Ste86d, SV86, Ste87c]. Expertise [Ste90a]. experts [Ste79b]. expertsystemen [NS86]. Exploiting [Ste90d, Ste90e].

Factors [SKMV02]. Fault [Ste89c]. Fifth [PBMW98]. finding [SBS91]. First [MW90, PVC^{+}99, Ste90g, SK01]. FORLI. [Ste76a]. formal [SV76]. Formalism [Ste76b]. Formation [Ste98e]. FORTRAN [Ste76a]. Fourth [MMM^{+}96]. Frame [Ste78, Ste91b, SS83]. frame-based [SS83]. Frames [Ste81]. Framework [Ste93a, Ste93d]. France [SDR^{+}00]. free [Ste75b]. French [Ste02d]. Function [Ste89c]. Function-Fault [Ste89c]. Functionality [Ste90f, Ste94d]. further [Ste94f]. future [NKR^{+}85, SSv94b].

Games [Ste98d, SK98b, SM99, Ste01b, Ste02b, Ste02c, SV97b]. Geheugens [Ste86b]. General [Ste79a]. Generation [JS99, Ste85d, Ste85e, Ste87d, SV89, Ste86d, SV86, Ste87c]. Germany [IPSK04]. Göteborg [PVC^{+}99]. Grammar [Ste00a, Ste02a]. grammars [Ste75a, SV76]. Grounded [Ste96e, Ste97g, SK99c, Ste01e, SK02, Ste98c]. Grounding [SV97b, Ste02b]. Guest [Ste95a].

hardware [Ste91a]. Heads [Ste99d, Ste01g]. Hoegaarden [SSv94b]. hominidus [Ste95e]. hommes [Ste02d]. Homo [Ste95e]. human [Ste02e]. Humanoid [Ste01f]. Hybrid [KKS95].

ihre [Ste00c]. II [dBHS87]. image [PVC^{+}99]. Information [GW95, Ste76c]. Innovation [SK98b]. integrated [Ste91a].

Intelligence [IPSK04, SC85, Ste87a, Ste88a, Ste88b, SS91, Ste94a, SB95, Ste95f, Ste95h, Ste95c, Ste96d, Ste03, dBHS87]. intelligens [Ste95e]. Intelligent [NKR^{+}85, Ste95k, Ste95a, Ste95j]. Intelligenz [Ste95c, Ste95d, Ste96a]. International [C^{+}94, IPSK04, MMM^{+}96, MW90, MBF^{+}00, PBMW98, SDR^{+}00]. Internetfluegeln [Ste00b]. Introduction [Ste84a, Ste86d]. Invited [Ste86a]. Issues [SSV96, Ste91a].

jeux [SV97a]. July [IPSK04, dBHS87].

Kennissystemen [Ste92]. Knowledge [GSV93, Ste85c, Ste87b, SL92, Ste93b, Ste93c, SL93, SM94, SSV94a, Ste84b, SSv94b, Ste94f]. Kognitive [Ste00c]. KRS [Ste84b, Ste85a]. kunstcreatie [Ste86f]. künstlichem [Ste95e].

langage [SV97a, SK00]. Language [Ste78, Ste97b, Ste97f, Ste98d, Ste98e, SK98b, Ste99b, Ste99c, Ste00d, Ste00e, Ste01b, Ste01e, Ste02b, Ste02c, Ste76c, Ste77, Ste79a, SV97b, Ste00f, SK01, Ste02f]. Languages [Ste86a, Ste75b]. Lausanne [FMN99]. Learning [SV86, SV89, SdB95, Ste97b, SK99b, Ste01e, Ste01f, vW97, SK01, Ste02e, Ste02f, dJS03]. Leben [Ste95e]. Leeds [SS91]. leur [KS00]. Level [SM94, Ste98e]. lexicale [SK99a]. Lexicon [SK98a]. Lexicons [Ste96c]. Life [Ste94a, SB95, ST95, FMN99]. line [Ste94d]. Linguistic [Ste98a]. linguistics [Ste76c]. Linguistique [Ste01c]. Liquid [KKS95]. Lisp [Ste90g, Ste86e, SCF^{+}86]. list [Ste76a].

Machine [vW97]. make [Ste94f].

Management [Ste93b, Ste93c]. manual [Ste76a]. Maschinen [Ste95e]. Massive [Ste87a]. Mathematical [Ste93d, Ste94g]. May [PVC^{+}99]. Meaning [Ste87b, Ste96e, Ste98e, SKMV02, SK01].

Mechanism [Ste78, Ste97d]. Memories [Ste98d, Ste86c]. Message [ADS^{+}87]. Methodology [Ste01d]. Mieux [Ste02d]. Mirror [Ste00e]. Model [Ste89c]. modeled [Ste79b]. Modeling [KKS95, Ste97f]. modular [Ste77]. monde [KS00]. Multi [Ste98b]. Multi-Agent [Ste98b]. Mythos [GW95].

Naming [SM99]. Natural [Ste78, Ste76c, Ste79a]. nature [SDR^{+}00]. Neurons [Ste00e]. nonlinear [SBS91].

Object [Ste83a, Ste84b, Ste83b, Ste84c]. Object-Oriented [Ste83a, Ste84b, Ste83b, Ste84c]. Objects [Ste94b]. office [NKR^{+}85]. OLB [Ste76a]. on-line [Ste94d]. Ontologies [Ste98b]. Ontwikkelingen [NS86]. ORBIT [Ste83b, Ste84c]. Organisation [Ste98e]. Organising [Ste97e]. Organization [Ste87e, Ste99c, Ste90b, Ste90c]. Organizing [Ste95i]. Oriented [Ste83a, Ste83b, Ste84b, Ste84c, Ste94c]. Origin [Ste98a]. Origine [SK00]. Origins [Ste95h, Ste96d, Ste97g, Ste97f, Ste98b, Ste98e, Ste00e, SKMV02, Ste98c, Ste02a].

package [Ste76a]. Paper [Ste86a]. papers [SSV94a, IPSK04]. Parallel [SDR^{+}00]. Parallelism [Ste87a]. Paris [SDR^{+}00]. Parsing [Ste75b]. passing [ADS^{+}87]. Path [SBS91]. people [Ste02d]. Perceptual [Ste97a]. Perceptually [Ste96e]. perspective [PSFSS89, PSFSS91, Ste89a]. perspectives [NKR^{+}85]. phonology [SO00]. physical [SSV96]. platforms [Ste91a]. population [SK99a]. posed [Ste94h]. Position [SSV94a]. possible [Ste95g]. PPSN [SDR^{+}00]. Practical [Ste90g]. Prague [vW97]. problem [SDR^{+}00, Ste89a]. Proceedings [Ste76c, Ste90g, C^{+}94, FMN99, MMM^{+}96, MW90, MBF^{+}00, PBMW98, PVC^{+}99, SDR^{+}00, vW97, SS91, SSv94b]. Process [SL92]. Processes [BKS00]. processing [PVC^{+}99, Ste76a, Ste79a, SS83]. Producing [Ste76c]. Programmeren [Ste86e]. Programming [Ste86a, Ste83b]. Progress [SC85, Ste94f]. properties [SV76, Ste79a]. puzzle [Ste00f].

reagieren [Ste00c]. Real [ST95]. realisation [SSV96]. Reasoning [Ste79b]. Recognition [Ste91b]. regular [Ste75b]. related [SV76]. Representation [Ste83a, Ste85c, Ste87b, Ste84b, Ste03]. Representations [Ste90e, Ste95f, Ste01e, Ste90d]. Republic [vW97]. Requirements [Ste85c]. research [Ste94h]. Reusability [Ste93a]. Revised [IPSK04]. Robo [Ste95e]. robot [Ste01a]. Roboter [Ste00c]. Robotic [Ste94d, Ste97g, Ste00a, SSV96, SV97b, Ste98c]. robotiques [SV97a, SK00]. Robots [Ste93d, SdB95, Ste01b, Ste01f, KS00, Ste95j]. Role [Ste93a, Ste01e]. Roots [Ste94a]. route [SB95].

sapiens [Ste95e]. Science [Ste97c, van95]. Second [Ste85d, Ste85e, Ste87c, Ste87d, SV89, Ste86d, SV86]. Selected [IPSK04]. Selection [Ste87e, JS99]. selectionist [Ste97d]. Self [Ste87e, Ste95i, Ste97e, Ste98e, Ste99c, Ste90b, Ste90c]. Self-Organisation [Ste98e]. Self-Organising [Ste97e]. Self-Organization [Ste87e, Ste99c, Ste90b, Ste90c]. Self-Organizing [Ste95i]. semantic [Ste76c]. Semantics [SK99c, SK02, Ste02c]. sémantique [SK99a]. Seminar [IPSK04]. Semiotic [SK99b]. Sense [Ste88d]. sensorielles [KS00]. sensory [JS99]. September [FMN99, SDR^{+}00, SSv94b]. session [Ste91a]. seventh [dBHS87]. Sharing [Ste97a]. Shift [Ste78]. sich [Ste00c]. signal [PVC^{+}99]. Simulation [C^{+}94, MMM^{+}96, MW90, MBF^{+}00, PBMW98, SS91]. simulations [Ste02a]. Situated [SB95, Ste99a, SK99c]. situés [SK99a]. Sixth [MBF^{+}00]. Social [Ste01f, Ste02e, Ste02f, SK01]. Society [SS91, Ste79b]. solving [SDR^{+}00, Ste89a]. Some [SS83]. Source [SK98b]. Spatial [Ste95i]. Spatially [SM99]. Spontaneous [SK98a, Ste99c]. Sprache [Ste00c]. standard [SCF^{+}86]. state [NKR^{+}85]. Steps [Ste88d]. Stochasticity [SK98b]. Structural [Ste98d]. Study [SS91, Ste94c]. style [Ste84c]. Subfield [Ste84a]. Sweden [PVC^{+}99]. Switzerland [FMN99]. Symbols [Ste02b]. syntactic [SS83, SO00]. Syntax [Ste97g, Ste98c]. Synthesising [Ste98e]. Synthetic [Ste97f]. System [Ste00d]. Systemen [Ste95c]. Systems [GSV93, Ste76b, Ste84a, Ste85c, Ste85d, Ste85e, Ste87d, Ste88c, Ste89b, SV89, Ste94e, SM94, Ste94g, Ste95b, Ste95h, Ste98b, ADS^{+}87, Ste75b, Ste79a, Ste84c, Ste86c, Ste86d, SV86, Ste87c, Ste91a].

Talking [Ste99d, Ste01g]. Technology [Ste95k, Ste95a]. telecommunications [PVC^{+}99]. Teleportation [Ste99a, Ste00c]. their [Ste75a, SV76]. Theory [Ste90f, Ste00e, Ste77]. Third [C^{+}94]. today [van95]. tool [Ste84c]. trends [van95].

U.K [dBHS87]. Umwelt [Ste00c]. understand [Ste02d]. University [SS91]. Ursprünge [Ste95c]. Use [Ste91b]. user [Ste76a]. Using [Ste98e, Ste84b].

Verbal [Ste01f]. VI [SDR^{+}00]. Vie [Ste01c]. view [Ste83b]. Visual [BSV98, SBV98]. Visually [Ste97g, Ste98c]. Vocabularies [Ste97e]. Vocabulary [Ste95i].

waves [SBS91]. Webb [Ste01d]. Welcome [GW95]. Wired [GW95]. Word [SK99c, SK02, SKMV02]. Word-Meaning [SKMV02]. words [SK01]. Workshop [SSv94b, SSV94a]. Workshops [PVC^{+}99]. workstation [NKR^{+}85]. World [GW95, ST95].

Zoekprocessen [Ste86f]. Zukunft [Ste95d, Ste96a].

References

Attardi:1987:MPA

[ADS^{+}87] G. Attardi, S. Diomedi, M. Simi, P. Maes, L. Steels, K. Van Marcke, C. Benoit, and C. Benoit. Message passing architectures and description systems. In Directorate General XIII, Telecommunications, Information Industries & Innovation, editor, ES-PRIT ’86: results and achievements, pages 73–86. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1987. ISBN 0-444-70160-5. LCCN A75.5 .E84 1986.

Birk:2000:EBP

[BKS00] A. Birk, H. Kenn, and L. Steels. Efficient behavioral processes. In Meyer et al. [MBF^{+}00], pages 391–398. ISBN 0-262-63200-4. ISSN 1089-4365. LCCN QL751.65.S55 I58 2000.

Belpaeme:1998:CAV

[BSV98] Tony Belpaeme, Luc Steels, and Joris Van Looveren. The construction and acquisition of visual categories. Lecture Notes in Computer Science, 1545:1–12, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/1545/15450001.htm;
http://link.springer-ny.com/
link/service/series/0558/papers/
1545/15450001.pdf.

Cliff:1994:AAP

[C^{+}94] Dave Cliff et al., editors. From animals to animats 3: proceedings of the Third International Conference on Simulation of Adaptive Behavior. MIT Press, MIT Press, 1994. ISBN 0-262-53122-4. LCCN QL751.65.S55 I58 1994.

duBoulay:1987:AAI

[dBHS87] Ben du Boulay, David Hogg, and Luc Steels, editors. Advances in artificial intelligence, II: seventh European Conference on Artificial Intelligence, ECAI-86, Brighton, U.K., July 20–25, 1986, volume 2. North-Holland, Amsterdam, The Netherlands, 1987. ISBN 0-444-70279-2. LCCN Q334.E97 1986.

deJong:2003:DLA

[dJS03] Edwin D. de Jong and Luc Steels.

A distributed learning algorithm for communication development. Complex Systems, 14(4):315–334, 2003. ISSN 0891-2513.

Floreano:1999:AAL

[FMN99] Dario Floreano, Francesco Mondada, and J. D. (Jean-Daniel) Nicoud, editors. Advances in artificial life: 5th European Conference, ECAL’99, Lausanne, Switzerland, September, 1999; proceedings, volume 1674 of Lecture Notes in Computer Science.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-66452-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .E87 1999.

Geldof:1993:CBK

[GSV93] S. Geldof, L. Steels, and W. Van de Velde. Commet for building knowledge systems. Communication and Cognition — Artificial Intelligence (CC-AI), 10(3):195– 218, 1993. ISSN 0378-0880.

Gerbel:1995:MIW

[GW95] Karl Gerbel and Peter Wiebel, editors. Mythos Information: Welcome to the Wired World: Ars Electronica 95. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-211-82709-9. LCCN QA76.9.C66 A77 1995.

Iida:2004:EAI

[IPSK04] Fumiya Iida, Rolf Pfeifer, Luc Steels, and Yasuo Kuniyoshi, editors. Embodied Artificial Intelligence: International Seminar, Dagstuhl Castle, Germany, July 7–11, 2003: Revised Selected Papers, volume 3139 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNCSD9. ISBN 3-540-22484-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q335 .E43 2004.

URL http://link.springer-
ny.com/link/service/series/

0558/tocs/t3139.htm;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=3139;   http://
www.springerlink.com/openurl.
asp?genre=volume&id=doi:10.
1007/b99075.

Jong:1999:GSS

[JS99] E. D. De Jong and L. Steels. Generation and selection of sensory channels. In Poli et al. [PVC^{+}99], pages 90–100. ISBN 3-540-65837-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5102.9 .E98 1999.

Kuyper:1995:HAM

[KKS95] J. De Kuyper, D. Keymeulen, and L. Steels. A hybrid architecture for modeling liquid behavior. In B. Chandrasekaran, Janice Glasgow, and N. Hari Narayanan, editors, Diagrammatic Reasoning: Cognitive and Computational Perspectives, pages 731– 751. AAAI Press, Menlo Park, CA, USA, 1995. ISBN 0-262-57112-9. LCCN QA76.9.A96 D53 1995.

Kaplan:2000:CRC

[KS00] F. Kaplan and L. Steels. Comment les robots construisent leur monde: Expériences sur la convergence des catégories sensorielles. In J-L. Dessalles, editor, Journée ARC Evolution et Cognition, pages 13– 18. ENST, Paris, France, 2000.

URL http://arti.vub.ac.be/
steels/arc2000.ps.

Landsbergen:1976:ANL

[LSB76] S. P. J. Landsbergen, L. Steels, and H. Boley. ???? In L. Steels, H. Boley, and S. P. J. Landsbergen, editors, Advances in natural language processing: Preprints of a workshop held at the University of Antwerp, 27 Oct. 1976, page 168. Universitaire Instelling Antwerpen, Wilrijk, Belgium, 1976.

Meyer:2000:AAP

[MBF^{+}00] Jean-Arcady Meyer, A. Berthoz, D. Floreano, H. L. Roitblat, and S. W. Wilson, editors. From animals to animats 6: proceedings of the Sixth International Conference on Simulation of Adaptive Behavior, Complex adaptive systems. MIT Press, MIT Press, 2000. ISBN 0-262-63200-4. ISSN 1089-4365. LCCN QL751.65.S55 I58 2000.

Maes:1996:AAP

[MMM^{+}96] Pattie Maes, M. J. Mataric, J.- A. Meyer, J. Pollack, and S. W. Wilson, editors. From animals to animats 4: proceedings of the Fourth International Conference on Simulation of Adaptive Behavior, Complex adaptive systems. MIT Press, MIT Press, 1996. ISBN 0-262-63178-4. ISSN 1089-4365. LCCN QL751.65.S55 I58 1996.

Meyer:1990:AAP

[MW90] Jean-Arcady Meyer and Stewart W. Wilson, editors. From animals to animats: proceedings of the First International Conference on Simulation of Adaptive Behavior. MIT Press, MIT Press, 1990. ISBN 0-262-63138-5. LCCN QL751.65.S55 I58 1990.

Naffah:1985:IWO

[NKR^{+}85] N. Naffah, G. Kempen, J. Rohmer, L. Steels, D. Tsichritzis, and G. White. Intelligent workstation in the office: state of the art and future perspectives. In J. Roukens and J. F. Renuart, editors, ESPRIT ’84: status report of ongoing work, pages 365– 378. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1985. ISBN 0-444-87740-1. LCCN QA75.5 .E84 1984.

Nijholt:1986:OE

[NS86] A. Nijholt and L. Steels, editors. Ontwikkelingen in expertsystemen. Academic Service, Den Haag, The Netherlands, 1986. ISBN 90-6233-202-1. iv + ch1(44) + ch2(14) + ch3(19) + ch4(18) + ch5(24) + ch6(25) + ch7(20) + ch8(18) pp.

Pfeifer:1998:AAP

[PBMW98] Rolf Pfeifer, B. Blumberg, J.-A.

Meyer, and S. W. Wilson, editors. From animals to animats 5: proceedings of the Fifth International Conference on Simulation of Adaptive Behavior, Complex adaptive systems. MIT Press, MIT Press, 1998. ISBN 0-262-66144-6. ISSN 1089-4365. LCCN QL751.65.S55 I58 1998.

Pfeifer:1989:CP

[PSFSS89] R. Pfeifer, Z. Schreter, F. Fogelman-Soulié, and L. Steels, editors. Connectionism in perspective. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1989. ISBN 0-444-88061-5. LCCN Q334 .C7 1989. Papers from an international conference held at the University of Zurich, Switzerland, 11–13 Oct. 1989.

Pfeifer:1991:CP

[PSFSS91] R. Pfeifer, Z. Schreter, F. Fogelman-Soulié, and L. Steels. Connectionism in perspective. Neurocomputing, 2(3):135, 1991. CODEN NR-CGEO. ISSN 0925-2312 (print), 1872-8286 (electronic).

Poli:1999:EIA

[PVC^{+}99] Riccardo Poli, H.-M. Voigt, S. Cagnoni, D. Corne, G. Smith, and T. Fogarty, editors. Evolutionary image analysis, signal processing, and telecommunications: First European Workshops, EvoIASP ’99 and EuroEcTel ’99, Göteborg, Sweden, May 26–27, 1999: proceedings, volume 1596 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-65837-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5102.9 .E98 1999.

Steels:1995:ALRa

[SB95] Luc Steels and Rodney Brooks, editors. The ‘Artificial Life’ route to ‘Artificial Intelligence’. Building Embodied, Situated Agents. Lawrence Erlbaum Associates, New Haven, CT, USA, 1995. ISBN 0-8058-1519-8 (paperback), 0-8058-1518-X (hardcover). ix + 288 pp. LCCN Q335 .A789 1995.

Sepulchre:1991:PFN

[SBS91] J. A. Sepulchre, A. Babloyantz, and L. Steels. Path finding with nonlinear waves. In Teuvo Kohonen et al., editors, Artificial neural networks: proceedings of the 1991 International Conference on Artificial Neural Networks (ICANN-91), Espoo, Finland, 24–28 June 1991, volume 2, pages 1265–1268. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1991. ISBN 0-444-89178-1. LCCN QA76.87.I56 1991. Two volumes.

Steels:1998:CAV

[SBV98] L. Steels, T. Belpaeme, and J. Van Looveren. The construction and acquisition of visual categories. In A. Birk and J. Demiris, editors, Learning robots: 6th European Workshop, EWLR ’96, Brighton, England, August 1996: proceedings, volume 1545 of Lecture Notes in Computer Science, pages 1–12. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-65480-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.I58A97

1998. URL http://arti.vub.
ac.be/steels/ewlr97.ps.

Steels:1985:PAI

[SC85] Luc Steels and J. A. Campbell, editors. Progress in Artificial Intelligence. Ellis Horwood, Chichester, UK, 1985. ISBN 0-85312-830-8, 0-470-20171-1 (Halsted Press). LCCN Q335 .P76 1985. A selection of papers and updates of papers from the 1982 European Conference on Artificial Intelligence in Orsay, France.

Stoyan:1986:TLS

[SCF^{+}86] H. Stoyan, J. Chailloux, J. Fitch, T. Krumnack, E. Neidl, J. Padget, G. Attardi, T. Christaller, J. Dalton, M. Devin, B. Lang, R. Lopez de Mantaras, E. Papon, S. Pope, C. Quiennec, and L. Steels. Towards a LISP standard. In du Boulay et al. [dBHS87], pages 46–52. ISBN 0-444-70279-2. LCCN Q334.E97 1986.

Steels:1995:LBB

[SdB95] L. Steels and B. de Boer. Learning basic behaviours on autonomous robots. In Proceedings of the 5th Belgian-Dutch Conference on Machine Learning, BENELEARN-95, page ?? ????, ????, 1995.

Schoenauer:2000:PPS

[SDR^{+}00] Marc Schoenauer, K. Deb, G. Rudolph, X. Yao, E. Lutton, J. J. Merelo, and H.-P. Schwefel, editors. Parallel problem solving from nature — PPSN VI; 6th International Conference, Paris, France, September 2000; proceedings, volume 1917 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-41056-2 (softcover). LCCN QA267.A1 L43 no.1917.

Steels:1998:SLC

[SK98a] L. Steels and F. Kaplan. Spontaneous lexicon change. In COLING-ACL ’98: 36th Annual Meeting of the Association for Computational Linguistics and 17th International Conference on Computational Linguistics, Aug. 10–14, 1998, Montreal, Quebec, Canada: proceedings of the conference, volume 2, pages 1243–1250. Morgan Kaufmann Publishers, San Francisco, CA, USA, 1998. ISBN 1-55860-581-9. LCCN P98 .A887 1998.

URL http://arti.vub.ac.be/
steels/acl98.ps.

Steels:1998:SSI

[SK98b] L. Steels and F. Kaplan. Stochasticity as a source of innovation in language games. In Christoph Adami, R. Belew, H. Kitano, and C. Taylor, editors, Artificial life VI: proceedings of the Sixth International Conference on Artificial Life, Complex adaptive systems, pages 368–376. MIT Press, MIT Press, 1998. ISBN 0-262-51099-5. ISSN 1089-4365. LCCN QH324.2 .I556 1998.

URL http://arti.vub.ac.be/
steels/alife98.ps. Conference

held at UCLA, June 27-29, 1998.

Steels:1999:ASL

[SK99a] L. Steels and F. Kaplan. Amorçage d’une sémantique lexicale dans une population d’agents autonomes, ancrés et situés. In P. Amsili, editor, 6e conférence annuelle sur le traitement automatique des langues naturelles, TALN’99, pages 393–398. ????, ????, 1999. URL http://arti.

                                        vub.ac.be/steels/taln99.ps.

Steels:1999:CLS

[SK99b] L. Steels and F. Kaplan. Collective learning and semiotic dynamics. In Floreano et al. [FMN99], pages 679–688. ISBN 3-540-66452-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .E87

                                        1999. URL http://arti.vub.
                                        ac.be/steels/collearn.ps.

Steels:1999:SGW

[SK99c] L. Steels and F. Kaplan. Situated grounded word semantics. In Thomas Dean, editor, IJCAI-99: proceedings of the Sixteenth International Joint Conference on Artificial Intelligence, Stockholm, Sweden, July 31–August 6, 1999, volume 2, pages 862–867. Morgan Kaufmann Publishers, San Francisco, CA, USA, 1999. ISBN 1-55860-613-0. LCCN Q334 .I7846 1999.

                                        URL http://arti.vub.ac.be/
                                        steels/ijcai99.ps. Two vol-

umes.

Steels:2000:OED

[SK00] L. Steels and F. Kaplan. Origine et évolution du langage: expériences robotiques. Revue du Palais de la Découverte, 278:63– 67, 2000.

Steels:2001:AFW

[SK01] L. Steels and F. Kaplan. AIBO’s first words. the social learning of language and meaning. Evolution of Communication, 4(1): ??, 2001. ISSN 1387-5337.

URL http://arti.vub.ac.be/
steels/aibo.ps. In press.

Steels:2002:BGW

[SK02] L. Steels and F. Kaplan. Bootstrapping grounded word semantics. In T. Briscoe, editor, Linguistic evolution through language acquisition: Formal and Computational Models, page ?? Cambridge University Press, Cambridge, UK, 2002. URL http://

arti.vub.ac.be/steels/ground.

ps. In press.

Steels:2002:CFO

[SKMV02] L. Steels, F. Kaplan, A. McIntyre, and J. Van Looveren. Crucial factors in the origins of word-meaning. In A. Wray, editor, The Transition to Language, pages 214–217. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002.

URL http://arti.vub.ac.be/
steels/crucialfac.ps.

Steels:1992:EKE

[SL92] Luc Steels and Brice Lepape, editors. Enhancing the Knowledge Engineering Process: Contributions from ESPRIT. Elsevier Science Publishers, Amsterdam, The Netherlands, 1992. ISBN 0-444-89893-X. vi + 273 pp. LCCN T58.5 .E54 1992.

Steels:1993:KEE

[SL93] L. Steels and B. Lepape. Knowledge engineering in ESPRIT. IEEE Intelligent Systems, 8(4):4– 10, 1993. ISSN 1094-7167.

Steels:1994:KLE

[SM94] Luc Steels and John McDermott, editors. The Knowledge Level in Expert Systems. Conversations and Commentary, volume 10 of Perspectives in artificial intelligence. Academic Press, San Diego, CA, USA, 1994. ISBN 0-12-664145-5. x + 274 pp. LCCN QA76.76.E95 S84 1994.

Steels:1999:SDN

[SM99] L. Steels and A. McIntyre. Spatially distributed naming games. Advances in Complex Systems, 1 (4):301–323, 1999. ISSN 0219-

                                        5259. URL http://arti.vub.
                                        ac.be/steels/spatial.ps.

Steels:2000:CES

[SO00] L. Steels and P.-Y. Oudeyer. The cultural evolution of syntactic constraints in phonology. In Mark A. Bedeau, J. S. McCaskill, N. H. Packard, and S. Rasmussen, editors, Artificial life VII: proceedings of the seventh International Conference on Artificial Life, pages 382–394. MIT Press, MIT Press, 2000. ISBN 0-262-52290-X. LCCN QH324.2 .I556

                                        2000. URL http://arti.vub.
                                        ac.be/steels/alife7.ps.

Steels:1983:SEF

[SS83] L. Steels and K. De Smedt. Some examples of frame-based syntactic processing. In Fr. Daems and L. Goossens, editors, Een spyeghel voor G. Jo Steenbergen, pages 293–305. Acco, Leuven, Belgium, 1983.

Steels:1991:APE

[SS91] Luc Steels and Barbara Smith, editors. AISB91: proceedings of the eighth Conference of the Society for the Study of Artificial Intelligence and Simulation of Behaviour, 16–19 April 1991, University of Leeds. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-19671-4 (New York), 3-540-19671-4 (Berlin). LCCN Q334 .S6315 1991.

Steels:1994:PPE

[SSV94a] L. Steels, G. Schreiber, and W. Van de Velde, editors. Position papers of the 8th European Knowledge Acquisition Workshop EKAW ‘94. Vrije Universiteit Brussel, Artificial Intelligence Laboratory, Technical Report 94-2, Brussels, Belgium, 1994. ISBN ???? LCCN ????

Steels:1994:FKA

[SSv94b] Luc Steels, Guus Schreiber, and Walter van de Velde, editors. A future for knowledge acquisition: 8th European Knowledge Acquisition Workshop, EKAW ’94, Hoegaarden, Belgium, September 26– 29, 1994: proceedings, volume 867 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 0-387-58487-0. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 E36 1994.

Steels:1996:IPR

[SSV96] L. Steels, P. Stuer, and D. Vereertbrugghen. Issues in the physical realisation of autonomous robotic agents. Unpublished notes., 1996.

                                        URL http://arti.vub.ac.be/
                                        steels/sab96.ps.

Steels:1995:ALRb

[ST95] L. Steels and M. Tokoro. Artificial life and real world computing. In Computer science today: recent trends and developments [van95], pages 15–28. CODEN LNCSD9. ISBN 3-540-60105-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76. .C57326 1995.

Steels:1975:CGT

[Ste75a] L. Steels. Completion grammars and their applications. Technical report, Universitaire Instelling Antwerpen, Wilrijk, Belgium, 1975. 63 pp.

Steels:1975:PSR

[Ste75b] L. Steels. Parsing systems for regular and context-free languages. Technical report, Universitaire Instelling Antwerpen, Wilrijk, Belgium, 1975. 44 pp.

Steels:1976:FOP

[Ste76a] L. Steels. The FORLI. OLB package for list processing in FOR-TRAN 4: a user’s manual. Technical report, Universitaire Instelling Antwerpen, Wilrijk, Belgium, 1976. 35 pp.

Steels:1976:FCS

[Ste76b] L. Steels. A formalism for case systems. In Proceedings of the Summer Conference on Artificial Intelligence and Simulation of Behaviour, AISB (ECAI’76), pages 279–290. ????, ????, 1976.

Steels:1976:PNL

[Ste76c] L. Steels. Producing natural language from semantic information: Proceedings of the computational linguistics conference. Technical report, Universitaire Instelling Antwerpen, Wilrijk, Belgium, 1976. 31 pp.

Steels:1977:AMT

[Ste77] L. Steels. Aspects of a modular theory of language, volume 50 of Proefschriften. Universitaire Instelling Antwerpen, Wilrijk, Belgium, 1977. 189 + 167 + 233 pp. Ph.D. Thesis.

Steels:1978:FSM

[Ste78] L. Steels. The frame shift mechanism in natural language. In D. H. Sleeman, editor, Proceedings of the 4th European Conference on Artificial Intelligence, AISB/GI (ECAI’78), pages 325– 329. Leeds University, Leeds, UK,

Steels:1979:GPN

[Ste79a] L. Steels. General properties of natural language processing systems. In Robert Trappl et al., editors, Progress in cybernetics and systems research, volume V, pages 504–510. Hemisphere Pub. Corp., New York, NY, USA, 1979. ISBN 0-470-88475-4. LCCN Q300 .P75. 11 volumes of papers presented at a symposium organized by the Austrian Society for Cybernetic Studies.

Steels:1979:RMS

[Ste79b] L. Steels. Reasoning modeled as a society of communicating experts. Technical Report AI-TR-542, MIT AI Lab, MIT Press, 1979. 154 pp.

Steels:1981:F

[Ste81] L. Steels. Frames. In Burghard B. Rieger, editor, Empirical Semantics: a collection of new approaches in the field, volume 12– 13 of Quantitative Linguistics, pages 192–235. Studienverlag Dr. N. Brockmeyer, Bochum, Germany, 1981. ISBN 3-88339-221-9 (vol. 1), 3-88339-222-7 (vol. 2). LCCN P325 .E42 1981.

Steels:1982:CC

[Ste82] L. Steels. Constraints as consultants. In Proceedings of the 5th European Conference on Artificial Intelligence, ECAI’82, pages 75–78. ????, ????, 1982.

Steels:1983:DCO

[Ste83a] L. Steels. Descriptions as constraints in object-oriented representation. In A. Bundy, editor, IJCAI-83 : proceedings of the Eighth International Joint Conference on Artificial Intelligence, 8–12 August 1983, Karlsruhe, West Germany, pages 395–397. William Kaufmann, San Francisco, CA, 1983. ISBN 0-86576-064-0. LCCN Q334 .I57 1983.

Steels:1983:OAV

[Ste83b] L. Steels. ORBIT: an applicative view of object-oriented programming. In Pierpaolo Degano and Erik Sandewall, editors, Integrated interactive computing systems: proceedings of the European Conference on Integrated Interactive Computing Systems, ECICS 82 Stresa, Italy, 1–3 September 1982, pages 193–205. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1983. ISBN 0-444-86595-0. LCCN QA76.9.I58 E96 1982.

Steels:1984:IES

[Ste84a] L. Steels. Introduction to expert systems subfield. In Tim O’Shea, editor, Advances in artificial intelligence: proceedings of the Sixth European Conference on Artificial Intelligence, ECAI-84, Pisa, Italy, September 5–7, 1984, page 1. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1984. ISBN 0-444-87611-1. LCCN Q334.E97 1984.

Steels:1984:OOK

[Ste84b] L Steels. Object-oriented knowledge representation using KRS. In Tim O’Shea, editor, Advances in artificial intelligence: proceedings of the Sixth European Conference on Artificial Intelligence, ECAI-84, Pisa, Italy, September 5–7, 1984, pages 333– 336. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1984. ISBN 0-444-87611-1. LCCN Q334.E97 1984.

Steels:1984:OTB

[Ste84c] L. Steels. ORBIT: a tool for building AI systems in an objectoriented style. In Alick Elithorn and Ranan Banerji, editors, Artificial and Human Intelligence: edited review papers presented at the International NATO Symposium on Artificial and Human Intelligence, Lyon, France, October 1981, pages 131–140. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1984. ISBN 0-444-86545-4. LCCN Q334 .I575 1981.

Steels:1985:ACK

[Ste85a] L. Steels. Algorithmic concepts in KRS. In NGI-SION 1985. Informatica Symposium: Uitdaging aan de informatica, pages 531– 542. Nederlands Genootschap voor Informatica. Stichting Informatica Onderzoek Nederland, Amsterdam, The Netherlands, 1985. ISBN 90-70621-17-7.

Steels:1985:CC

[Ste85b] L. Steels. Constraints as consultants. In Luc Steels and J. A. Campbell, editors, Progress in Artificial Intelligence, pages 146– 165. Ellis Horwood, Chichester, UK, 1985. ISBN 0-85312-830-8. LCCN Q335 .P76 1985. A selection of papers and updates of papers from the 1982 European Conference on Artificial Intelligence in Orsay, France.

Steels:1985:DRK

[Ste85c] L. Steels. Design requirements for knowledge representation systems. In Joachim Laubsch, editor, GWAI-84: 8th German Workshop on Artificial Intelligence, Wingst/Stade, October 8–12, 1984, volume 103 of Informatik-Fachberichte: Subreihe Kunstliche Intelligenz, pages 1–19. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-15669-0 (New York), 3-540-15669-0 (Berlin). LCCN Q334 .G47 1984.

Steels:1985:SGEa

[Ste85d] L. Steels. Second generation expert systems. Informatie, 27 (11):966–971, 1985. CODEN IN-FTCR. ISSN 0019-9907.

Steels:1985:SGEb

[Ste85e] Luc Steels. Second generation expert systems. Future Generation Computer Systems, 1(4):213–221, June 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Steels:1986:APL

[Ste86a] L. Steels. AI and programming languages (invited paper). In H.-J. Kugler, editor, Information processing 86: proceedings of the IFIP 10th World Computer Congress, Dublin, Ireland, September 1–5, 1986, volume 10 of IFIP congress series, pages 105–110. North-Holland, Amsterdam, The Netherlands, 1986. ISBN 0-444-70077-3. LCCN QA75.5.I3785 1986.

Steels:1986:CG

[Ste86b] L. Steels. Collectieve geheugens. In A. Nijholt and L. Steels, editors, Ontwikkelingen in expertsystemen, pages 8.1–8.19. Academic Service, Den Haag, The Netherlands, 1986.

Steels:1986:ESC

[Ste86c] L. Steels. From expert systems to Community Memories. In Thomas Bernold, editor, Expert systems and knowledge engineering: essential elements of advanced information technology: proceedings of the Technology Assessment and Management Conference of the Gottlieb Duttweiler Institute, Ruschlikon, Zurlich, Switzerland, 25–26, April 1985, pages 17–29. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1986. ISBN 0-444-70057-9. LCCN QA76.76.E95 T43 1985.

Steels:1986:ISG

[Ste86d] L. Steels. Introduction to second generation expert systems. In L. C. Aiello, editor, Nuove frontiere dell’informatica: i sistemi esperti, pages 131–144. Accademia Nazionale dei Lincei, Rome, Italy, 1986.

Steels:1986:PL

[Ste86e] L. Steels. Programmeren in LISP.

Academic Service, Den Haag, The Netherlands, 1986. ISBN 90-6233-096-7 (1983), 90-6233-229-3 (1986). iii + ch1(22) + ch2(44) + ch3(34) + ch4(24) + ch5(33) + ch6(13) + ch7(6) + bib(5) pp.

Steels:1986:ZK

[Ste86f]  L. Steels.  Zoekprocessen in

kunstcreatie. In R. Hagoort and P. Maessen, editors, Geest, computer, kunst, pages 286–296. Stichting Grafiet, Utrecht, The Netherlands, 1986. ISBN 90-70559-07-2.

Steels:1987:MPA

[Ste87a] L. Steels. Massive parallelism for artificial intelligence. Microprocessing and Microprogramming, 21(1–5):17–19, 1987. CO-DEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic).

Steels:1987:ERM

[Ste87b] L. Steels. Meaning in knowledge representation. In Pattie Maes and Daniele Nardi, editors, Meta-Level Architectures and Reflection, pages 51–59. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1987. ISBN 0-444-70343-8. LCCN QA76.7 .M48 1988. Selection of the papers presented at the Workshop Meta-level Architectures and Reflection, held in Alghero, Italy, during October 27– 30, 1986 and sponsored by the Cost-13 Project No. 21.

Steels:1987:SGEa

[Ste87c] L. Steels. Second generation expert systems. In Proceedings of the SEAS Spring Meeting 1987: Systems Architecture. 2 volumes, pages 101–109. SHARE European Association, Nijmegen, The Netherlands, 1987.

Steels:1987:SGEb

[Ste87d] L. Steels. Second generation expert systems. In M. A. Bramer, editor, Research and development in expert systems III: proceedings of the Sixth Technical Conference of the British Computer Society Specialist Group on Expert Systems, Metropole Hotel, Brighton, 16–18 December 1986, pages 175–183. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34154-X. LCCN QA76.76.E95 B76 1986.

Steels:1987:SOT

[Ste87e]  L. Steels.  Self-organization

through selection. In M. Caudill and C. Butler, editors, IEEE First International Conference on Neural Networks, Sheraton Harbor Island East, San Diego, California, June 21–24, 1987, volume 2, pages 55–62. SOS Printing, Piscataway, NJ, USA, 1987. LCCN QA76.87 .I3437 1987; QP363.3 .I58 1987; QT 34 I22 1987i. Four volumes. IEEE catalog number 87 THO191-7.

Steels:1988:AICa

[Ste88a] L. Steels. Artificial intelligence and complex dynamics. In Third International Symposium on Knowledge Engineering, pages 164–168. Rank Xerox, Madrid, Spain, 1988.

Steels:1988:AICb

[Ste88b] L. Steels. Artificial intelligence and complex dynamics. In Mario Tokoro, Yuichiro Anzai, and Akinori Yonezawa, editors, Concepts and characteristics of knowledgebased systems: selected and reviewed papers from the IFIP TC 10/WG 10.1 workshop, Mount Fuji, Japan, 9–12 November, 1987, pages 369–404. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1988. ISBN 0-444-88328-2. LCCN QA76.76.E95 C664 1989.

Steels:1988:DES

[Ste88c] L. Steels. The deepening of expert systems. AI Communications, 1 (1):9–17, 1988. CODEN ACM-MEE. ISSN 0921-7126.

Steels:1988:STC

[Ste88d] L. Steels. Steps towards common sense. In Yves Kodratoff, editor, ECAI 88: proceedings of the 8th European Conference on Artificial Intelligence: August 1–5, 1988, pages 49–54. Pitman Publishing, London, UK, 1988. ISBN 0-273-08798-3. LCCN Q334 .E97 1988.

Steels:1989:CPS

[Ste89a] L Steels. Connectionist problem solving: an AI perspective. In Pfeifer et al. [PSFSS89], pages 215–228. ISBN 0-444-88061-5. LCCN Q334 .C7 1989. Papers from an international conference held at the University of Zurich, Switzerland, 11–13 Oct. 1989.

Steels:1989:DES

[Ste89b] L. Steels. The deepening of expert systems. In R. Huber et al., editors, Artificial Intelligence in Scientific Computation: Towards second generation systems, IMACS’89, pages 189–196. J.C. Baltzer AG, Scientific Publishing Co., Basel, Switzerland, 1989.

Steels:1989:DFF

[Ste89c] L. Steels. Diagnosis with a function-fault model. Applied Artificial Intelligence, 3(2-3):213– 237, 1989.

Steels:1990:CE

[Ste90a] L. Steels. Components of expertise. The AI Magazine, 11(2): 28–49, 1990. CODEN AIMAEK. ISSN 0738-4602.

Steels:1990:CBDa

[Ste90b] L. Steels. Cooperation between distributed agents through selforganization. In Y. Demazeau and J.-P. Müller, editors, IROS ’90: towards a new frontier of applications: proceedings / IEEE International Workshop on Intelligent Robots and Systems ’90, July 3–6, 1990, Mechanical Engineering Research Laboratory, Hitachi Ltd., Tsuchiura, Ibaraki, Japan, pages 8–14 suppl. IEEE, New York, NY, USA, 1990. ISBN ???? LCCN TJ210.3 .I443 1990. Two volumes. IEEE catalog number: 90TH0332-7.

Steels:1990:CBDb

[Ste90c] L. Steels. Cooperation between distributed agents through selforganization. In Yves Demazeau and Jean-Pierre Müller, editors, Decentralized A.I.: proceedings of the First European Workshop on Modelling Autonomous Agents in a Multi-Agent World, Cambridge, England, August 16–18, 1989, pages 175–196. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1990. ISBN 0-444-88705-9. LCCN Q334.E99 1989.

Steels:1990:EARa

[Ste90d] L. Steels. Exploiting analogical representations. Robotics and Autonomous Systems, 6(1–2):71–88, 1990. CODEN RASOEJ. ISSN 0921-8890. URL http://www.

sciencedirect.com/science/
journal/09218890.

Steels:1990:EARb

[Ste90e] L. Steels. Exploiting analogical representations. In Pattie Maes, editor, Designing Autonomous Agents: Theory and Practice from Biology to Engineering and Back, pages 71–88. MIT Press, MIT Press, 1990. ISBN 0-262-63135-0. LCCN Q325 .D47 1991.

Steels:1990:TTE

[Ste90f] L. Steels. Towards a theory of emergent functionality. In Meyer and Wilson [MW90], pages 451– 461. ISBN 0-262-63138-5. LCCN QL751.65.S55 I58 1990.

Steels:1990:EEC

[Ste90g] Luc Steels, editor. Europal 90: Proceedings of the First European Conference on the Practical Application of Lisp, Churchill College, Cambridge. Europal, Dorking, UK, 1990. ISBN ???? LCCN ????

Steels:1991:DSH

[Ste91a] L. Steels. Discussion session on hardware issues. the choice of development and delivery platforms for integrated systems. In Expert systems integration: proceedings of the BANKAI Workshop, Brussels, Belgium, 12–14 September 1990, pages 51–56. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1991. ISBN 0-444-88649-4. LCCN QA76.76.E95 B26 1990.

Steels:1991:EFR

[Ste91b] L. Steels. Emergent frame recognition and its use in artificial creatures. In John Mylopoulos and Ray Reiter, editors, IJCAI-91: proceedings of the Twelfth International Conference on Artificial Intelligence, Darling Harbour, Sydney, Australia, 24-30 August 1991, volume 2, pages 1219–1224. Morgan Kaufmann Publishers, San Francisco, CA, USA, 1991. ISBN 1-55860-160-0. LCCN Q334 .I57 1991. Two volumes.

Steels:1992:K

[Ste92] L. Steels. Kennissystemen. Addison-Wesley, Reading, MA, USA, 1992. ISBN 90-6789-327-7. xviii + 554 pp.

Steels:1993:CFR

[Ste93a]  L. Steels.  The componen-

tial framework and its role in reusability. In Jean-Marc David, Jean-Paul Krivine, and Reid Simmons, editors, Second Generation Expert Systems, pages 273– 298. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-56192-7 (New York), 3-540-56192-7 (Berlin). LCCN QA76.76.E95 S43 1993.

Steels:1993:CKMa

[Ste93b] L. Steels. Corporate knowledge management. In Jose Cuena, editor, Knowledge oriented software design: extended papers from the IFIP TC 12 Workshop on Artificial Intelligence from the Information Processing Perspective, AIFIPP ’92, Madrid, Spain, 14–15 September, 1992, volume A-27 of IFIP Transactions: A. Computer science and technology, pages 91– 116. North-Holland, Amsterdam, The Netherlands, 1993. ISBN 0-444-81541-4. ISSN 0926-5473. LCCN QA76.76.D47 I35 1992.

Steels:1993:CKMb

[Ste93c] L. Steels. Corporate knowledge management. In J.-P. Barthès, editor, Proceedings of the 1st International Symposium on the Management of Industrial and Corporate Knowledge, IS-MICK’93, pages 9–30. ????, ????, 1993.

Steels:1993:MFA

[Ste93d] L. Steels. A mathematical framework for autonomous robots. In Patrick Brezillon and Vadim Stefanuk, editors, Proceedings of the East-West Conference on Artificial Intelligence: EWAIC ’93, September 7–9, 1993, Moscow, Russia, pages 333–334. ????, ????, 1993. ISBN ???? LCCN Q334 .E18 1993.

Steels:1994:ALR

[Ste94a] L. Steels. The artificial life roots of artificial intelligence. Artificial Life, 1(1–2):89–125, 1994. CO-DEN ARLIEY. ISSN 1064-5462, 1530-9185. URL http://arti.

vub.ac.be/steels/alife.ps.

Steels:1994:BO

[Ste94b] L. Steels. Beyond objects. In Mario Tokoro and Remo Pareschi, editors, Object-oriented programming: 8th European Conference, ECOOP ’94, Bologna, Italy, July 4–8, 1994: proceedings, volume 821 of Lecture Notes in Computer Science, pages 1– 11. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CO-DEN LNCSD9. ISBN 0-387-58202-9 (New York), 3-540-58202-9 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .E95 1994.

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/0821/08210001.htm;

 http://link.springer-ny.com/
link/service/series/0558/papers/
0821/08210001.pdf.

Steels:1994:CSB

[Ste94c] L. Steels. A case study in the behavior-oriented design of autonomous agents. In Dave Cliff, Ph. Husbands, J.-A. Meyer, and S. W. Wilson, editors, From Animals to Animats 3. Proceedings of the Third International Conference on Simulation of Adaptive Behavior, SAB’94, Complex Adaptive Systems, pages 445–452. MIT Press, MIT Press, 1994. ISBN 0-262-53122-4. LCCN QL751.65.S55 I58 1994.

URL http://arti.vub.ac.be/
steels/sab94.ps.

Steels:1994:EFR

[Ste94d] L. Steels. Emergent functionality in robotic agents through on-line evolution. In Rodney Brooks and Pattie Maes, editors, rtificial life IV: proceedings of the Fourth International Workshop on the Synthesis and Simulation of Living Systems, pages 8–14. MIT Press, MIT Press, 1994. ISBN 0-262-52190-3. LCCN QH324.2 .A75

1994. URL http://arti.vub.
ac.be/steels/alife4.ps.

Steels:1994:EAB

[Ste94e] L. Steels. Equilibrium analysis of behavior systems. In A. G. Cohn, editor, ECAI 94, 11th European Conference on Artificial Intelligence, August 8–12, 1994, Amsterdam, the Netherlands: proceedings, pages 714–718. John Wiley and Sons, Ltd., Chichester, UK, 1994. ISBN 0-471-95069-6. LCCN Q334.E97 1994.

Steels:1994:HCW

[Ste94f] L. Steels. How can we make further progress in knowledge acquisition? In Proceedings of the Japanese Knowledge Acquisition Workshop, pages 65–71. ????, ????, 1994.

Steels:1994:MAB

[Ste94g] L. Steels. Mathematical analysis of behavior systems. In P. Gaussier and J.-D. Nicoud, editors, Proceedings: From Perception to Action Conference, September 7–9, 1994, Lausanne, Switzerland, pages 88– 95. IEEE, New York, NY, USA, 1994. ISBN 0-8186-6482-7 (case), 0-8186-6481-9 (microfiche). LCCN TJ210.3 .F76

                                        1994. URL http://arti.vub.
                                        ac.be/steels/perarc.ps.

Steels:1994:NCP

[Ste94h] L. Steels. New challenges posed by autonomous agents research. In T. Lengauer, editor, Computer Science: State of the Art and Future Directions, page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994.

Steels:1995:BTIb

[Ste95a] L. Steels. The biology and technology of intelligent autonomous agents (guest editorial). Robotics and Autonomous Systems, 15(1– 2):1–2, 1995. CODEN RASOEJ. ISSN 0921-8890.

Steels:1995:BAA

[Ste95b] L. Steels. Building agents with autonomous behavior systems. In Steels and Brooks [SB95], pages 83–119. ISBN 0-8058-1519-8 (paperback), 0-8058-1518-X (hardcover). LCCN Q335 .A789 1995.

URL http://arti.vub.ac.be/
steels/corsendonck.ps.

Steels:1995:UIA

[Ste95c]  L. Steels.  Die Ursprünge

von Intelligenz bei Artificial-Intelligence-Systemen. In Gerbel and Wiebel [GW95], pages 308– 312. ISBN 3-211-82709-9. LCCN QA76.9.C66 A77 1995.

Steels:1995:ZI

[Ste95d] L. Steels. Die Zukunft der Intelligenz. In S. Bollmann, editor, Kursbuch Neue Medien: Trends in Wirtschaft und Politik, Wissenschaft und Kultur, pages 321–333. Bollmann Verlag, Mannheim, Germany, 1995.

Steels:1995:HCS

[Ste95e] L. Steels. Homo cyber-sapiens oder Robo hominidus intelligens: Maschinen erwachen zu künstlichem Leben. In Ch. Maar, E. Pöppel, and T. Christaller, editors, Die Technik auf dem Weg zur Seele: Proceedings of the Burda Symposium on Brain-Computer Interfaces, pages 327– 344. Rowohlt Taschenbuch Verslag, Hamburg, Germany, 1995.

URL http://arti.vub.ac.be/

                                        steels/burda.ps. German ver-

sion of The Homo Cyber Sapiens, the Robot Homonidus Intelligens, and the ‘artificial life’ approach to artificial intelligence.

Steels:1995:IDR

[Ste95f] L. Steels. Intelligence — dynamics and representations. In Steels [Ste95k], pages 72–89. ISBN 3-540-59052-8. LCCN TJ217.5 .B56

                                        1995. URL http://arti.vub.
                                        ac.be/steels/trento.ps. Pro-

ceedings of the NATO Advanced Study Institute on the Biology and Technology of Intelligent Autonomous Agents, held in Castel Ivano, Trento, Italy, March 1–12, 1993.

Steels:1995:ACP

[Ste95g] L. Steels. Is artificial consciousness possible? In G. Trautteur, editor, Consciousness: Distinction and Reflection. Bibliopolis, Naples, Italy, 1995.

Steels:1995:OIA

[Ste95h] L. Steels. The origins of intelligence in artificial systems. In Gerbel and Wiebel [GW95], pages 308–312. ISBN 3-211-82709-9. LCCN QA76.9.C66 A77 1995.

Steels:1995:SOS

[Ste95i] L. Steels. A self-organizing spatial vocabulary. Artificial Life, 2 (3):319–332, 1995. CODEN AR-LIEY. ISSN 1064-5462, 1530-

                                        9185. URL http://arti.vub.
                                        ac.be/steels/space.ps.

Steels:1995:WRI

[Ste95j] L. Steels. When are robots intelligent autonomous agents? Robotics and Autonomous Systems, 15(1–2):3–9, 1995. CO-DEN RASOEJ. ISSN 0921-

8890. URL http://arti.vub.
ac.be/steels/autsys.ps.

Steels:1995:BTIa

[Ste95k] Luc Steels, editor. The Biology and Technology of Intelligent Autonomous Agents, volume 144 of NATO ASI Series: Series F, Computer and Systems Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-59052-8. LCCN TJ217.5 .B56 1995. Proceedings of the NATO Advanced Study Institute on the Biology and Technology of Intelligent Autonomous Agents, held in Castel Ivano, Trento, Italy, March 1–12, 1993.

Steels:1996:ZI

[Ste96a] L. Steels. Die Zukunft der Intelligenz. Kunstforum International, 133:84–89, 1996. ISSN 0177-3674.

Steels:1996:DC

[Ste96b] L. Steels. Discovering the competitors. Adaptive Behavior, 4(2):173–199, 1996. CODEN ADBEEA. ISSN 1059-7123.

URL http://arti.vub.ac.be/
steels/para.ps.

Steels:1996:EAL

[Ste96c]  L. Steels.  Emergent adap-

tive lexicons. In Maes et al.

[MMM^{+}96], pages 562–567. ISBN 0-262-63178-4. ISSN 1089-4365. LCCN QL751.65.S55 I58 1996.

URL http://arti.vub.ac.be/
steels/adaptlex.ps.

Steels:1996:OI

[Ste96d] L. Steels. The origins of intelligence. In The Future of Science Has Begun: Approaches to Artificial Life and Artificial Intelligence. Proceedings of the Fourth International Conference, pages 75–88. Fondazione Carlo Erba, Milano, Italy, 1996.

URL http://arti.vub.ac.be/
steels/origin.ps.

Steels:1996:PGM

[Ste96e] L. Steels. Perceptually grounded meaning creation. In M. Tokoro, editor, Proceedings of the Second International Conference on Multi-Agent Systems, ICMAS-96, pages 338–344. AAAI Press, Menlo Park, CA, USA, 1996.

URL http://arti.vub.ac.be/
steels/meaning.ps.

Steels:1997:CSP

[Ste97a] L. Steels. Constructing and sharing perceptual distinctions. In van Someren and Widmer [vW97], pages 4–13. CO-DEN LNCSD9. ISBN 3-540-62858-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .E85 1997.

URL http://arti.vub.ac.be/
steels/ecml.ps.

Steels:1997:LLL

[Ste97b] L. Steels. Language learning and language contact. In W. Daelemans, A. Van den Bosch, and A. Weijters, editors, Workshop Notes of the ECML/MLnet Workshop on Empirical Learning of Natural Language Processing Tasks, ECML-97, pages 11–24. Prague, Czech Republic, 1997.

URL http://arti.vub.ac.be/
steels/contact.ps.

Steels:1997:SCA

[Ste97c] L. Steels. Science communication in the age of cyberspace. In The Future of Science has begun: The communication of science to the public. Science and the media. Proceedings of the Fifth International Conference, pages 173–180. Fondazione Carlo Erba, Milano, Italy, 1997.

Steels:1997:SMA

[Ste97d] L. Steels. A selectionist mechanism for autonomous behavior acquisition. Robotics and Autonomous Systems, 20(2–):117– 131, 1997. CODEN RASOEJ. ISSN 0921-8890.

Steels:1997:SOV

[Ste97e] L. Steels. Self-organising vocabularies. In Christopher G. Langton and Katsunori Shimohara, editors, Proceedings of the 5th International Workshop on Artificial Life: Synthesis and Simulation of Living Systems, ALIFE-96, pages 179–184. MIT Press, MIT Press, 1997. ISBN 262621118. LCCN QH324.2

                                        .I56 1996. URL http://arti.
                                        vub.ac.be/steels/alife96.ps.

Workshop held at the Nara-Ken New Public Hall, Nara, Japan, May 16th to May 18th, 1996.

Steels:1997:SML

[Ste97f] L. Steels. The synthetic modeling of language origins. Evolution of Communication, 1(1): 1–34, 1997. ISSN 1387-5337.

                                        URL http://arti.vub.ac.be/
                                        steels/coe.ps.

Steels:1997:OSV

[Ste97g] Luc Steels. The origins of syntax in visually grounded robotic agents. In Martha E. Pollack, editor, IJCAI-97: proceedings of the Fifteenth International Joint Conference on Artificial Intelligence, Nagoya, Japan, August 23–29, 1997, volume 2, pages 1632–1641. Morgan Kaufmann Publishers, San Francisco, CA, USA, 1997. ISBN 1-55860-480-4. LCCN Q334 .I57 1997.

                                        URL http://arti.vub.ac.be/
                                        steels/orig-syn.ps. Two vol-

umes.

Steels:1998:OLC

[Ste98a] L. Steels. The origin of linguistic categories. In The Evolution of Language Conference: Selected papers from 2nd International Conference on the Evolution of Language, page ?? ????, ????,

                                        1998. URL http://arti.vub.
                                        ac.be/steels/orig-cat.ps.

Steels:1998:OOC

[Ste98b] L. Steels. The origins of ontologies and communication conventions in multi-agent systems. Autonomous Agents and Multi-Agent Systems, 1(2):169– 194, 1998. ISSN 1387-2532.

URL http://arti.vub.ac.be/
steels/orig-ont.ps.

Steels:1998:OSV

[Ste98c] L. Steels. The origins of syntax in visually grounded robotic agents. Artificial Intelligence, 103(1-2):133–156, 1998. CO-DEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (elec-

tronic). URL http://arti.vub.
ac.be/steels/orig-syn.ps.

Steels:1998:SCC

[Ste98d] L. Steels. Structural coupling of cognitive memories through adaptive language games. In Pfeifer et al. [PBMW98], pages 263–269. ISBN 0-262-66144-6. ISSN 1089-4365. LCCN QL751.65.S55 I58 1998.

Steels:1998:SOL

[Ste98e]  L. Steels.  Synthesising the

origins of language and meaning using co-evolution, selforganisation and level formation. In James R. Hurford, Michael Studdert-Kennedy, and Chris Knight, editors, Approaches to the Evolution of Language: social and cognitive bases, pages 384– 404. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-63049-5, 0-521-63964-6 (paperback). LCCN P116 .A68 1998.

URL http://arti.vub.ac.be/
steels/edin.ps.

Steels:1999:CTS

[Ste99a] L. Steels. Cognitive teleportation and situated embodiment. In Floreano et al. [FMN99], pages 7–8. ISBN 3-540-66452-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QH324.2 .E87 1999.

Steels:1999:HLB

[Ste99b] L. Steels. How language bootstraps cognition. In I. Wachsmutt and B. Jung, editors, KogWis99. Proceedings der 4. Fachtagung der Gesellschaft für Kognitionswissenschaft, pages 1–3. Infix, Braunschweig, Germany, 1999.

Steels:1999:SSO

[Ste99c]  L. Steels.  The spontaneous

self-organization of an adaptive language. In K. Furukawa, D. Michie, and S. Muggleton, editors, Intelligent agents, volume 15 of Machine Intelligence, pages 205–224. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-853867-7 (hardcover). LCCN QA76.9.M35 I53 1992.

                                        URL http://arti.vub.ac.be/
                                        steels/mi15.ps.

Steels:1999:THE

[Ste99d] L. Steels. The Talking Heads Experiment. Laboratorium, Antwerpen, Belgium, 1999. 238 pp. Limited Pre-edition.

Steels:2000:EGC

[Ste00a] L. Steels. The emergence of grammar in communicating autonomous robotic agents. In Werner Horn, editor, ECAI 2000: 14th European Conference on Artificial Intelligence, August 20– 25, 2000, Berlin, Germany: including Prestigious Applications of Intelligent Systems (PAIS-2000): proceedings, volume 54 of Frontiers in artificial intelligence and applications, pages 764–769. IOS Press, Amsterdam, The Netherlands, 2000. ISBN 1-58603-013-2 (IOS Press), 4-274-90388-5 (Ohmsha). ISSN 0922-

6389. URL http://arti.vub.
ac.be/steels/ecai2000.ps.

Steels:2000:EIG

[Ste00b] L. Steels. Engeln mit Internetfluegeln. In Hans-Joachim Schellnhuber et al., editors, Die Gegenwart der Zukunft: die Serie der Süddeutschen Zeitung über unsere Welt im neuen Jahrhundert, volume 367 of Wagenbach Taschenbuch, pages 90–98. Verlag Klaus Wagenbach, Berlin, Germany, 2000. ISBN 3-8031-2367-4.

URL http://www.benecke.com/
gegenwart.html. German ver-

sion of Digital Angels.

Steels:2000:KRT

[Ste00c] L. Steels. Kognitive Roboter und Teleportation — Artefacte reagieren auf ihre Umwelt und erfinden sich eine Sprache. In Ch Maar, H. U. Obrist, and E. Pöppel, editors, Weltwissen Wissenswelt: Das globale Netz von Text und Bild, pages 175–184. DuMont Verlag, Cologne, Germany, 2000.

Steels:2000:LCA

[Ste00d] L. Steels. Language as a complex adaptive system. In Schoenauer et al. [SDR^{+}00], pages 17–26. ISBN 3-540-41056-2 (softcover). LCCN QA267.A1 L43 no.1917.

                                        URL http://arti.vub.ac.be/
                                        steels/ppsn2000.ps.

Steels:2000:MNA

[Ste00e] L. Steels. Mirror neurons and the action theory of language origins. Unpublished notes., 2000.

                                        URL http://arti.vub.ac.be/
                                        steels/neurons.ps.

Steels:2000:PLE

[Ste00f] L. Steels. The puzzle of language evolution. Kognitionswissenschaft, 8(4):143–150, 2000. CODEN KOGNEB. ISSN 0938-

                                        7986. URL http://arti.vub.
                                        ac.be/steels/puzzle.ps.

Steels:2001:CRC

[Ste01a] L. Steels. Can a robot be conscious? In press., 2001.

Steels:2001:LGA

[Ste01b] L. Steels. Language games for autonomous robots. IEEE Intelligent Systems, 16(5):16– 22, 2001. ISSN 1094-7167.

                                        URL http://arti.vub.ac.be/
                                        steels/games.ps.

Steels:2001:LEV

[Ste01c] L. Steels. Linguistique evolutionnaire et vie artificielle. Vie Artificielle, 1(1):??, 2001.

Steels:2001:MAC

[Ste01d] L. Steels. The methodology of the artificial. commentary on Barbara Webb’s article. Behavioral and brain sciences, 24(6): ??, 2001. CODEN BBSCDH. ISSN 0140-525X (print), 1469-1825 (electronic). In press.

Steels:2001:RLL

[Ste01e] L. Steels. The role of language in learning grounded representations. In P. R. Cohen and T. Oates, editors, Learning Grounded Representations: Papers from 2001 AAAI Spring Symposium, number SS-01-05 in AAAI Spring Symposium Series Technical Reports, pages 80–88. AAAI Press, Menlo Park, CA, USA, 2001.

Steels:2001:SLV

[Ste01f] L. Steels. Social learning and verbal communication with humanoid robots. In Proceedings of the IEEE-RAS International Conference on Humanoid Robots, pages 335–342. IEEE, New York, NY, USA, 2001.

Steels:2001:THE

[Ste01g] L. Steels. The talking heads experiment. In H. U. Obrist and B. Vanderlinden, editors, Laboratorium, pages 413–419. DuMont Verlag, Cologne, Germany, 2001.

Steels:2002:CSO

[Ste02a] L. Steels. Computer simulations of the origins of case grammar. In press. To be presented at evolang2002 in Harvard, in preparation., 2002.

Steels:2002:GST

[Ste02b] L. Steels. Grounding symbols through evolutionary language games. In Angelo Cangelosi and Domenico Parisi, editors, Simulating the Evolution of Language, pages 211–226. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 1-85233-428-2. LCCN P116 .C36 2001.

Steels:2002:LGE

[Ste02c] L. Steels. Language games for emergent semantics. IEEE Intelligent Systems, 17(1):83–85, 2002. ISSN 1094-7167.

Steels:2002:MCH

[Ste02d] L. Steels. Mieux comprendre les hommes. . .. (French) [To better understand people…]. La Recherche: Les nouveaux robots (special issue), 350:72–76, 2002.

Steels:2002:SCL

[Ste02e] L. Steels. Social and cultural learning in the evolution of human communication. In D. K. Oller, U. Griebel, and K. Plunkett, editors, Evolution of Communication Systems: A Comparative Approach, page ?? MIT Press, MIT Press, 2002. In press.

Steels:2002:SLL

[Ste02f] L. Steels. Social learning and language acquisition. In D. McFarland, O. Holland, and C. Melhuish, editors, Social Robots, page ?? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. In press.

Steels:2003:IR

[Ste03] Luc Steels. Intelligence with representation. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 361(1811): 2381–2395, 2003. CODEN PTR-MAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic). Biologically inspired robotics (Bristol, 2002).

Steels:1976:FPC

[SV76] L. Steels and D. Vermeir. On the formal properties of completion grammars and their related automata. Technical report, Universitaire Instelling Antwerpen, Wilrijk, Belgium, 1976. 41 pp.

Steels:1986:LSG

[SV86] L. Steels and W. Van de Velde.

Learning in second generation expert systems. In Janusz S. Kowalik, editor, Knowledge-based problem solving, pages 270–295. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-516576-8. LCCN QA76.76.E95 K561 1986.

Steels:1989:LSG

[SV89] L. Steels and W. Van de Velde.

Learning in second generation expert systems. In Yves Kodratoff and Alan Hutchinson, editors, Machine and Human Learning. Advances in European Research, Michael Horwood series on artificial intelligence, pages 53–75. Kogan Page, London, UK, 1989. ISBN 1-85091-497-4. LCCN Q325.M317 1989b.

Steels:1997:AJL

[SV97a] L. Steels and P. Vogt. Ancrage de jeux de langage adaptatifs dans des agents robotiques. In A. Ali Cherif and J. Signorini, editors, Intelligence Artificielle et Complexité, pages 27–39. Université Paris 8, 1997. ISBN 2-9511847-0-0.

Steels:1997:GAL

[SV97b] L. Steels and P. Vogt. Grounding adaptive language games in robotic agents. In Phil Husbands and Inman Harvey, editors, Proceedings of the Fourth European Conference on Artificial Life, ECAL’97, Complex Adaptive Systems, pages 474–482. MIT Press, MIT Press, 1997. ISBN 0-262-58157-4. LCCN QH324.2

.E87 1997. URL http://arti.
vub.ac.be/steels/grounding.

ps. Conference held July 28-31, 1997 in Brighton, UK.

vanLeeuwen:1995:CST

[van95] J. van Leeuwen. Computer science today: recent trends and developments, volume 1000 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CODEN LNCSD9. ISBN 3-540-60105-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 641 pp. LCCN QA76. .C57326 1995.

Someren:1997:MLE

[vW97] Maarten W. van Someren and Gerhard Widmer, editors. Machine learning: ECML-97: 9th European Conference on Machine Learning, Prague, Czech Republic, April 23–25, 1997: proceedings, volume 1224 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNCSD9. ISBN 3-540-62858-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q325.5 .E85 1997.